# Field-Level Comparison and Robustness Analysis of Cosmological N-Body Simulations

### v1

@@ 1 Introduction
To perform inference, simulation-based inference (SBI) is an emerging tool to determine the cosmological parameters by training a neural network on cosmological simulations (see e.g. Villaescusa-Navarro et al., 2021a; Hahn et al., 2024; Ho et al., 2024; Tucci & Schmidt, 2024). While SBI can be applied to any summary of the data, such as the power spectrum, here we focus on field-level inference which seeks to extract all of the information from every pixel in the field (see e.g. Jasche & Wandelt, 2013; Seljak et al., 2017; Jasche & Lavaux, 2019; Schmidt et al., 2019, 2020; Nguyen et al., 2021; Kostić et al., 2023; Bayer et al., 2023, 2023; Lemos et al., 2024; Sharma et al., 2024; Nguyen et al., 2024; Andrews et al., 2024; Jia, 2024; Horowitz & Lukic, 2025; Parker et al., 2025).
@@
More broadly in astrophysics and machine learning, there has been much recent attention to the sensitivity of SBI methods to distribution shifts (Filipp et al., 2024; Swierc et al., 2024; Agarwal et al., 2024), with some attempts made to alleviate these issues using domain adaptation and generalization (Anau Montel & Weniger, 2022; Masserano et al., 2023). To understand the implications for cosmological N-body simulations, we quantify the distributional shifts between the outputs different cosmological simulations by performing a statistical out-of-distribution (OOD) analysis using PQMass (Lemos et al., 2024).
@@
  • While existing works comparing different N-body simulations consider a collection of standard summaries, such as the power spectrum and bispectrum (e.g. Schneider et al., 2016; Garrison et al., 2019; Grove et al., 2022), it is known that these do not fully describe the total information in the cosmic field. Thus, here we compare the simulations at the field level. In particular, we perform a statistical OOD analysis using PQMass (Lemos et al., 2024) to quantify the comparability of the density fields produced by different simulations.
@@ 2.1 Simulations
We run simulations with different seeds in a Latin Hypercube setup with parameter ranges $0.1\leq\Omega_{\rm m}\leq 0.5$ and $0.6\leq\sigma_{8}\leq 1.0$, as well as 1 simulation with matched seed and cosmology ($\Omega_{m}=0.3175$ and $\sigma_{8}=0.834$). For Gadget we generate 1000 simulation in the Latin Hypercube for training, while for the other codes we use $\sim 50$ simulations drawn from the same distribution of cosmological parameters for testing, except for Enzo and Gizmo for which we do no Latin Hypercube analysis. For the matched seed datasets we ran an extra set of simulations for each code with double the particle resolution. The Latin Hypercubes are used to quantify the precision and accuracy on parameter inference, while the matched seed and cosmology simulations are used to understand the robustness in a controlled setting.
@@
    Ramses. An AMR code based on the technique described in Teyssier (2002). The AMR framework is based on a graded octree, where cells are refined individually when the enclosed mass exceeds a certain multiple $n_{\rm max}$ of the particle mass. For our initial analysis we refine when a cell has $n_{\rm max}=8$ times the particle mass, and also investigate the sensitivity to this parameter using $n_{\rm max}=2$ (denoted ’Ramses2’) for one specific part of our analysis in 3.4. Mass deposition is done using the Cloud-In-Cell method. Poisson’s equation is solved level by level using Dirichlet boundary conditions from the coarser level and a Multigrid relaxation solver. Time integration is performed with the Verlet algorithm (also known as adaptive leapfrog). The minimum refinement level was set to $\ell_{\rm min}=8$, corresponding to 2563 base grid cells (one per particle on average), and the maximum level was $\ell_{\rm max}=15$, with a minimum cell size of about $1\,{\rm kpc}/h$. Note that the refinement strategy we have adopted here for the Enzo code corresponds to $n_{\rm max}=3$.
@@ 2.3 Neural Network for Parameter Inference
We use a CNN architecture which takes a $256^{2}$ input image and processes it through six convolutional blocks, each consisting of multiple convolutional layers, batch normalization, and LeakyReLU activations, progressively increasing the feature depth and downsampling spatial dimensions. The model uses circular padding to avoid edge artifacts. After flattening the output, two fully connected layers predict a 12-dimensional output, where the last six values are squared to enforce positivity. Dropout is applied for regularization. This architecture is optimized for learning hierarchical features from the input while minimizing overfitting.
@@
[…]
@@
To compare the field-level results of the network, and robustness to different N-body solvers, we train the network using Gadget simulations, and test using all of the different simulations. The data is split into three different sets: training, validation, and testing in the ratio $90\%:5\%:5\%$, such that each 2D map comes from a different original simulations box. During training, we perform data augmentation through rotations and flips to enforce invariance to these symmetries. We use the optuna package (Akiba et al., 2019) to perform hyperparameter optimization over the number of hidden layers, the dropout rate for the fully connected layers, the learning rate, and the weight decay. We consider various metrics of network performance, including the residual error $\epsilon$, the root mean square error (RMSE), the coefficient of determination $R^{2}$, and the $\chi^{2}$.
@@ 3.2 Field-Level Out-of-Distribution Analysis
Figure 3: Field-level OOD analysis. PQMass $\chi^{2}_{\rm PQM}$ distributions, comparing each simulation with respect to Gadget. Under the null hypothesis, and given our use of 50 comparison samples, when two sets of samples are from the same underlying distribution, $\chi^{2}_{\rm PQM}$ should follow a $\chi^{2}$ distribution with 49 degrees of freedom (black line). When comparing Gadget to Gadget, as a null test, $\chi^{2}_{\rm PQM}$ closely follows the $\chi^{2}$ distribution, implying Gadget is in distribution with itself. However, all other simulations tested against Gadget yield $\chi^{2}_{\rm PQM}$ distributions that deviate significantly from the expected $\chi^{2}$ behavior, indicating they are OOD relative to Gadget. Abacus, CUBEP3M, and PKDGrav are similarly OOD with a mean $\chi^{2}_{\rm PQM}$ of around 64, while Ramses and TNG are the most OOD with a $\chi^{2}_{\rm PQM}$ of around 77.
@@
Fig. 3 presents the distribution of $\chi^{2}_{\rm PQM}$ (eqn. (2.2)) obtained by comparing each simulation with respect to Gadget. As we use $n_{R}=50$ regions, the expected distribution is a $\chi^{2}$ distribution with 49 degrees of freedom, that is, a mean of 49 and a standard deviation of 9.89, depicted by the black line. As expected, the $\chi^{2}_{\rm PQM}$ from the Gadget null test is in agreement with the theoretical $\chi^{2}$ distribution, confirming internal consistency among Gadget samples (the small discrepancy in the standard deviation could be improved by using more samples). In contrast, all other simulations tested against Gadget yield $\chi^{2}_{\rm PQM}$ distributions that deviate significantly from the expected $\chi^{2}$ behavior, indicating that they are OOD relative to Gadget. We see that Abacus, CUBEP3M, and PKDGrav are similarly OOD with a mean $\chi^{2}_{\rm PQM}$ of around 64, while Ramses and TNG are the most OOD with a $\chi^{2}_{\rm PQM}$ of around 77.
@@ 3.3 Field-Level Inference
Figure 5: SBI results when training on Gadget simulations, and testing on all simulations with fixed seed and cosmology for $\Omega_{m}$ (top) and $\sigma_{8}$ (bottom). The 15 data points for each test correspond to 5 $5\,{\rm Mpc}/h$ slices that fit in the box sliced in all 3 directions. We perform tests at two particle resolutions, one where the network is tested at the same resolution as the training set, and another where the network is tested on simulations with twice the resolution (HR). When trained and tested at the same resolution, the agreement between non-AMR codes is overall good, while the AMR codes (Enzo and Ramses) show a bias, predicting a high $\Omega_{m}$ and a low $\sigma_{8}$. However, when testing the network on HR simulations we find a significant bias in all cases. This suggested that the model misspecification in AMR codes is due to differences in the effective resolution of the training and testing set, which particularly affects the AMR simulations as they have an adaptive resolution.
@@ 4 Conclusions
We showed that smoothing the fields enabled robust inference between simulations with different resolutions. While techniques such as domain adaptation seek a shared latent space across simulations, here we instead use physically motivated filtering to align the simulations by removing scales where differences are expected. It would be interesting future work to explore alternative methods of filtering the data to potentially more optimally ensure robust inference. For example, Lahiry et al. (2025) showed that applying cuts in the density to exclude the most extreme regions does not remove much cosmological information regarding $\Omega_{m}$ and $\sigma_{8}$. Moreover, while Shao et al. (2023) considered graph neural networks applied to halos from these simulations, it would be interesting to apply the CNN approach used here on maps of halos or galaxies to find the cuts on parameters (such as minimum halo mass) required to ensure robustness between different N-body simulations and halo finders.

### v2

@@ 1 Introduction
To perform inference, simulation-based inference (SBI) is an emerging tool to determine the cosmological parameters by training a neural network on cosmological simulations (see e.g. Villaescusa-Navarro et al., 2021a; Hahn et al., 2024; Ho et al., 2024; Tucci & Schmidt, 2024). While SBI can be applied to any summary of the data, such as the power spectrum, here we focus on field-level inference which seeks to extract all of the information from every pixel in the field (see e.g. Jasche & Wandelt, 2013; Seljak et al., 2017; Jasche & Lavaux, 2019; Schmidt et al., 2019, 2020; Nguyen et al., 2021; Jeffrey et al., 2021; Kostić et al., 2023; Bayer et al., 2023, 2023; Lemos et al., 2024; Sharma et al., 2024; Nguyen et al., 2024; Andrews et al., 2024; Jia, 2024; Horowitz & Lukic, 2025; Golshan & Bayer, 2025; Parker et al., 2025).
@@
More broadly in astrophysics and machine learning, there has been much recent attention to the sensitivity of SBI methods to distribution shifts (Filipp et al., 2024; Swierc et al., 2024; Agarwal et al., 2024) and adversarial attacks (Horowitz & Melchior, 2022), with some attempts made to alleviate these issues using domain adaptation and generalization (Anau Montel & Weniger, 2022; Masserano et al., 2023). To understand the implications for cosmological N-body simulations, we quantify the distributional shifts between the outputs different cosmological simulations by performing a statistical out-of-distribution (OOD) analysis using PQMass (Lemos et al., 2024).
@@
  • While existing works comparing different N-body simulations consider a collection of standard summaries, such as the power spectrum and bispectrum (e.g. Schneider et al., 2016; Garrison et al., 2019; Grove et al., 2022), or use wavelets to quantify the noise in simulations (Romeo et al., 2008), it is known that these do not fully unpack the total information in the cosmic field. Thus, here we compare the simulations at the field level. In particular, we perform a statistical OOD analysis using PQMass (Lemos et al., 2024) to quantify the comparability of the density fields produced by different simulations.
@@ 2.1 Simulations
We run simulations with different seeds in a Latin Hypercube setup with parameter ranges $0.1\leq\Omega_{\rm m}\leq 0.5$ and $0.6\leq\sigma_{8}\leq 1.0$, as well as 1 simulation with matched seed and cosmology ($\Omega_{m}=0.3175$ and $\sigma_{8}=0.834$). The other cosmological parameters are fixed at $\Omega_{b}=0.049$, $h=0.6711$, $n_{s}=0.9624$, $M_{\nu}=0$, and w=1w=-1. For Gadget we generate 1000 simulation in the Latin Hypercube for training, while for the other codes we use $\sim 50$ simulations drawn from the same distribution of cosmological parameters for testing, except for Enzo and Gizmo for which we do no Latin Hypercube analysis. For the matched seed datasets we ran an extra set of simulations for each code with double the particle resolution. The Latin Hypercubes are used to quantify the precision and accuracy on parameter inference, while the matched seed and cosmology simulations are used to understand the robustness in a controlled setting.
@@
    Ramses. An AMR code based on the technique described in Teyssier (2002). The AMR framework is based on a graded octree, where cells are refined individually when the enclosed mass exceeds a certain multiple $n_{\rm max}$ of the particle mass. For our initial analysis we refine when a cell has $n_{\rm max}=8$ times the particle mass, and also investigate the sensitivity to this parameter using $n_{\rm max}=2$ (denoted ‘Ramses2’) for one specific part of our analysis in 3.4. Mass deposition is done using the Cloud-In-Cell method. Poisson’s equation is solved level by level using Dirichlet boundary conditions from the coarser level and a Multigrid relaxation solver. Time integration is performed with the Verlet algorithm (also known as adaptive leapfrog). The minimum refinement level was set to $\ell_{\rm min}=8$, corresponding to 2563 base grid cells (one per particle on average), and the maximum level was $\ell_{\rm max}=15$, with a minimum cell size of about $1\,{\rm kpc}/h$. Note that the refinement strategy we have adopted here for the Enzo code corresponds to $n_{\rm max}=3$.
@@ 2.3 Neural Network for Parameter Inference
We use a CNN architecture which takes a $256^{2}$ input image and processes it through six convolutional blocks, each consisting of multiple convolutional layers, batch normalization, and LeakyReLU activations, progressively increasing the feature depth and downsampling spatial dimensions. The model uses circular padding to avoid edge artifacts. After flattening the output, two fully connected layers predict a 12-dimensional output (corresponding to the 6 parameter means and standard deviations), where the last six values are squared to enforce positivity. Dropout is applied for regularization. This architecture is optimized for learning hierarchical features from the input while minimizing overfitting. The architecture is described fully in the Appendix in Table 2.
@@
$\mu_{i}({X})=\int_{\theta_{i}}{p(\theta_{i}|\bm{X})\theta_{i}d\theta_{i}},$ (3)
$\sigma_{i}({X})=\int_{\theta_{i}}{p(\theta_{i}|\bm{X})(\theta_{i}-\mu_{i})^{2}d\theta_{i}},$ (4)
@@
To compare the field-level results of the network, and robustness to different N-body solvers, we train the network using Gadget simulations, and test using all of the different simulations. The data is split into three different sets: training, validation, and testing in the ratio $90\%:5\%:5\%$, such that each 2D map comes from a different original simulations box. During training, we perform data augmentation through rotations and flips to enforce invariance to these symmetries. We use the Adam optimizer with a cyclic learning rate scheduler with minimum learning rate of $10^{-9}$. We use the optuna package (Akiba et al., 2019) to perform hyperparameter optimization over the number of hidden layers $H$ (between 6 and 12), the dropout rate for the fully connected layers (between 0 and 0.9), the weight decay (between $10^{-8}$ and $10^{-1}$), and the maximum learning rate (between $10^{-5}$ and $10^{-3}$). We consider various metrics of network performance, including the residual error $\epsilon$, the root mean square error (RMSE), the coefficient of determination $R^{2}$, and the $\chi^{2}$.
@@ 3.2 Field-Level Out-of-Distribution Analysis
Figure 3: Field-level OOD analysis. PQMass $\chi^{2}_{\rm PQM}$ distributions, comparing each simulation with respect to Gadget. Under the null hypothesis, and given our use of 50 comparison samples, when two sets of samples are from the same underlying distribution, $\chi^{2}_{\rm PQM}$ should follow a $\chi^{2}$ distribution with 49 degrees of freedom (black line). When comparing Gadget to Gadget, as a null test, $\chi^{2}_{\rm PQM}$ closely follows the $\chi^{2}$ distribution, implying Gadget is in distribution with itself. However, all other simulations tested against Gadget yield $\chi^{2}_{\rm PQM}$ distributions that deviate significantly from the expected $\chi^{2}$ behavior, indicating they are OOD relative to Gadget. Abacus, CUBEP3M, and PKDGrav are similarly OOD with a mean $\chi^{2}_{\rm PQM}$ of between 61-65, while Ramses and TNG are the most OOD with a $\chi^{2}_{\rm PQM}$ of around 75.
@@
Fig. 3 presents the distribution of $\chi^{2}_{\rm PQM}$ (eqn. (2.2)) obtained by comparing each simulation with respect to Gadget. As we use $n_{R}=50$ regions, the expected distribution is a $\chi^{2}$ distribution with 49 degrees of freedom, that is, a mean of 49 and a standard deviation of 9.89, depicted by the black line. As expected, the $\chi^{2}_{\rm PQM}$ from the Gadget null test is in agreement with the theoretical $\chi^{2}$ distribution, confirming internal consistency among Gadget samples (the small discrepancy in the standard deviation could be improved by using more samples). In contrast, all other simulations tested against Gadget yield $\chi^{2}_{\rm PQM}$ distributions that deviate significantly from the expected $\chi^{2}$ behavior, indicating that they are OOD relative to Gadget. We see that Abacus, CUBEP3M, and PKDGrav are similarly OOD with a mean $\chi^{2}_{\rm PQM}$ of between 61-65, while Ramses and TNG are the most OOD with a $\chi^{2}_{\rm PQM}$ of around 75.
@@ 3.3 Field-Level Inference
Figure 5: SBI results when training on Gadget simulations, and testing on all simulations with fixed seed and cosmology for $\Omega_{m}$ (top) and $\sigma_{8}$ (bottom). The 15 data points for each test correspond to 5 $5\,{\rm Mpc}/h$ slices that fit in the box sliced in all 3 directions. We perform tests at two particle resolutions, one where the network is tested at the same resolution as the training set, and another where the network is tested on simulations with twice the resolution (HR). When trained and tested at the same resolution, the agreement between non-AMR codes is overall good, while the AMR codes (Enzo and Ramses) show a bias, predicting a high $\Omega_{m}$ and a low $\sigma_{8}$. However, when testing the network on HR simulations we find a significant bias in all cases. This suggests that the model misspecification in AMR codes is due to differences in the effective resolution of the training and testing set, which particularly affects the AMR simulations as they have an adaptive resolution.
@@ 4 Conclusions
We showed that smoothing the fields enabled robust inference between simulations with different resolutions. While techniques such as domain adaptation seek a shared latent space across simulations, here we instead use physically motivated filtering to align the simulations by removing scales where differences are expected. It would be interesting future work to explore alternative methods of filtering the data to potentially more optimally ensure robust inference. For example, Lahiry et al. (2025) showed that applying cuts in the density to exclude the most extreme regions does not remove much cosmological information regarding $\Omega_{m}$ and $\sigma_{8}$. It would also be fruitful to explore methods to perform robust inference by training using multiple datasets, such as Jo et al. (2025). Moreover, while Shao et al. (2023) considered graph neural networks applied to halos from these simulations, it would be interesting to apply the CNN approach used here on maps of halos or galaxies to find the cuts on parameters (such as minimum halo mass) required to ensure robustness between different N-body simulations and halo finders.